What is the frequency of a wave with a wavelength 12 meters travelling at 68 m/s?
a
5.67 Hz
b
0.71 Hz
c
4.12 hertz
d
3.56 hertz
e
1.44 Hz
Question 4 (1 point)
1
Expert's answer
2018-12-19T10:32:11-0500

frequency = 1/period or f=1/T; speed of the wave = wavelength/period ; v = λ/T ; v = λ * f ; f = v/λ = 68/12=5.6. Answer A) 5.67 Hz